# On tori periods of Weil representations of unitary groups

### v1

@@ Abstract.
We determine the restriction of Weil representations of unitary groups to maximal tori. In the local case, we show that the Weil representation contains a pair of compatible characters if and only if a root number condition holds. In the global case, we show that a tori period corresponding to a maximal anisotropic torus of the global theta lift of a character does not vanish if and only if the local condition is satisfied everywhere and a central value of an $L$-function does not vanish. Our proof makes use of the seesaw argument and of the well-known theta lifting results from $\operatorname{U}(1)$ to $\operatorname{U}(1)$. Our results are used in [1, 2] to construct Arthur packets for $G_{2}$.
@@ 1. Introduction
Branching problems are a fascinating topic in representation theory and in the theory of automorphic representations. The most famous examples are the Gan–Gross–Prasad conjectures [9, 10, 11] and their refinements [18, 16, 32], extending the original conjectures of Gross–Prasad [13, 14]. In this paper, we study the restriction of Weil representations of unitary groups to maximal tori. Our results are both local and global, and they are similar in nature to the Gan–Gross–Prasad conjectures. Let us describe the problems we concern. Let $F$ be a field with characteristic different than $2$ and let $K/F$ be a quadratic étale algebra with involution $x\mapsto x^{c}$, whose set of fixed points is $F$. Let $\mathrm{V}$ be a non-degenerate $n$-dimensional hermitian space over $K$, and let $\mathrm{W}$ be a non-degenerate one-dimensional skew-hermitian space over $K$. When $F$ is a local field, we consider the following branching problem: given a maximal torus $\mathrm{T}$ of $\operatorname{U}(\mathrm{V})$ and characters $\alpha\colon\mathrm{T}\to\mathbb{C}^{\times}$ and $\beta\colon\operatorname{U}(\mathrm{W})\to\mathbb{C}^{\times}$, we would like to investigate whether the restriction of the Weil representation of the metaplectic group $\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$ to $\mathrm{T}\times\operatorname{U}(\mathrm{W})$ contains the representation $\alpha\boxtimes\beta$ as a sub-quotient. Reformulating this using the theta correspondence, this is equivalent to asking whether the space $\operatorname{Hom}_{\mathrm{T}}\left(\Theta\left(\beta\right),\alpha\right)$ is non-zero, where $\Theta\left(\beta\right)$ is the big theta lift of $\beta$ from $\operatorname{U}(\mathrm{W})$ to $\operatorname{U}(\mathrm{V})$. Suppose that $F=\mathbf{F}$ is a number field and that $\mathbf{K}/\mathbf{F}$ is a quadratic field extension. For an algebraic group $G$, we write $\left[G\right]=G\left(\mathbf{F}\right)\backslash G\left(\mathbb{A}_{\mathbf{F}}\right)$. We consider the following branching problem: given a maximal torus $\mathrm{T}$ of $\operatorname{U}(\mathrm{V})$ and automorphic characters $\alpha\colon\left[\mathrm{T}\right]\to\mathbb{C}^{\times}$ and $\beta\colon\left[\operatorname{U}(\mathrm{W})\right]\to\mathbb{C}^{\times}$, we would like to investigate whether the $\alpha$-period of the global theta lift $\Theta\left(\beta\right)$ of $\beta$ from $\operatorname{U}(\mathrm{W})\left(\mathbb{A}_{\mathbf{F}}\right)$ to $\operatorname{U}(\mathrm{V})\left(\mathbb{A}_{\mathbf{F}}\right)$ is non-zero. That is, we are asking whether the assignment
@@
is identically zero on the space $\Theta\left(\beta\right)$. In order to avoid convergence issues, we will assume that the torus $\mathrm{T}$ is anisotropic in the global setting, so that the integrals in question converge absolutely. Notice that when $\mathrm{V}$ is one-dimensional, the theta lift of $\beta$ is either zero or a character, and our problems reduce to determining whether $\Theta\left(\beta\right)$ equals $\alpha$ or not. This problem, of determining the theta lift from $\operatorname{U}(1)$ to $\operatorname{U}(1)$, is well understood. It dates back to Moen [23], Rogawski [27], and Harris–Kudla–Sweet [15] in the non-archimedean local field case, to Paul [24] in the archimedean local field case, to Minguez [22], Fang–Sun–Xue [5] and Gan [7] in the split local case, and to Rogawski [27], Yang [35] and Yamana [34] in the global case. See also Section 9 of [8] and the last paragraph of Section 7 of [8]. Our technique for solving these problems in the general case, where $\dim\mathrm{V}$ is arbitrary, involves invoking a seesaw identity that reduces the problems to the well-known case discussed above. This idea has been used before by Gan and his collaborators, see for example [6, Sections 9 and 10], [7, Section 5] and [8, Section 10]. In order to state our results, we need a classification of maximal tori in $\operatorname{U}(\mathrm{V})$. In Section 2, we recall the classification given in [26]. Each maximal torus $\mathrm{T}\subset\operatorname{U}(\mathrm{V})$ corresponds to an étale algebra $E$ of degree $n$ over $F$ and an element $\lambda\in E^{\times}$, such that the space $\left(\mathrm{V}_{E,\lambda},\left\langle\cdot,\cdot\right\rangle_{\lambda}\right)$ is isomorphic to $\mathrm{V}$ as hermitian spaces, where $\mathrm{V}_{E,\lambda}=\operatorname{Res}_{L_{E}/K}L_{E}$, equipped with the hermitian product $\left\langle x,y\right\rangle_{\mathrm{V}_{E,\lambda}}=\operatorname{tr}_{L_{E}/K}{\left(\lambda xy^{c}\right)}$, where $L_{E}=K\otimes_{F}E$. In this case, the maximal torus $\mathrm{T}_{E,\lambda}$ is isomorphic to the norm one torus of $L_{E}$, that is,
@@
We now move to describe our main results. Suppose that $F$ is a local field. By choosing characters $\left(\chi_{\mathrm{V}},\chi_{\mathrm{W}}\right)$ of $K^{\times}$ we obtain a splitting $\tilde{\iota}\colon\operatorname{U}(\mathrm{V})\times\operatorname{U}(\mathrm{W})\to\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$ of the metaplectic group. Let $i^{\prime}_{\mathrm{W}}\colon\operatorname{U}(\mathrm{W})\to L_{F}^{1}$ be an admissible embedding, which in this case is an isomorphism. Suppose that $E$ is an étale algebra of degree $n$ over $F$. Let $\beta\colon L_{F}^{1}\to\mathbb{C}^{\times}$, and $\alpha\colon L_{E}^{1}\to\mathbb{C}^{\times}$ be characters. We give the following answer (Theorem 5.1) to the local problem discussed above.
@@
$\operatorname{Hom}_{i\left(L_{E}^{1}\right)}\left(\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right),\alpha\circ i^{-1}\right)$ (1)
@@
    Character compatibility: $\beta=\left(\chi_{\mathrm{W}}^{-n}\cdot\chi_{\mathrm{V}}\right)\circ j_{F}^{-1}\cdot\alpha_{\restriction_{K^{1}}}.$
@@
    Root number condition: $\omega_{L_{E}/E}\left(\lambda\right)=\varepsilon_{L_{E}/E}\left(\alpha_{L_{E}}\cdot\chi_{\mathrm{W}}^{-1}\circ N_{L_{E}/K},\psi,\delta\right).$
@@
Here, $\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$ is the big theta lift from $\operatorname{U}(\mathrm{W})$ to $\operatorname{U}(\mathrm{V})$, and $0\neq\delta\in K^{\times}$ is a trace zero element, such that $\delta=\operatorname{disc}\mathrm{W}\left(\bmod N_{K/F}\left(K^{\times}\right)\right)$. We refer the reader to Sections 2.2.2 and 3.3 for the definition of the vector of quadratic characters $\omega_{L_{E}/E}\left(\lambda\right)$, the vector of root numbers $\varepsilon_{L_{E}/E}\left(\alpha_{L_{E}}\cdot\chi_{\mathrm{W}}^{-1}\circ N_{L_{E}/K},\psi,\delta\right)$, and other notation appearing in the theorem. We also show that for characters $\alpha$ and $\beta$ satisfying the compatibility condition, there exists a unique non-degenerate hermitian space $\mathrm{H}$ of dimension $n$, up to isomorphism, and a unique admissible embedding $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{H})$, up to conjugation, such that the space (1) attached to $i$ is non-zero. More precisely, we show the following theorem (Corollary 5.3).
@@
$\sum_{\mathrm{H}\in\mathrm{Her}_{n}}\sum_{i\in\Sigma_{E,\mathrm{H}}}\dim_{\mathbb{C}}\operatorname{Hom}_{i\left(L_{E}^{1}\right)}\left(\Theta_{\mathrm{W},\mathrm{H}}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right),\alpha\circ i^{-1}\right)=\begin{dcases}1&\beta=\left(\chi_{\mathrm{W}}^{-n}\cdot\chi_{\mathrm{V}}\right)\circ j_{F}^{-1}\cdot\alpha_{\restriction_{K^{1}}},\\ 0&\text{otherwise.}\end{dcases}$
@@
[…]
$i$ runs over representatives of classes of $\Sigma_{E,\mathrm{H}}$, and $\Theta_{\mathrm{W},\mathrm{H}}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$ is the theta lift from $\operatorname{U}(\mathrm{W})$ to $\operatorname{U}(\mathrm{H})$, taken with respect to the splitting defined by a prescribed choice of characters $\left(\chi_{\mathrm{W}},\chi_{\mathrm{V}}\right)$.
@@
$\mathcal{V}_{\mathrm{W},E}\left(\beta\right)=\bigcup_{\mathrm{H}\in\mathrm{Her}_{n}}\left\{\left(\Theta_{\mathrm{W},\mathrm{H}}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right),\alpha\circ i^{-1}\right)\mid i\in\Sigma_{E,\mathrm{H}}\right\},$
is analogous to the local Vogan $L$-packet appearing in the Gan–Gross–Prasad conjectures, consisting of irreducible representations of the group and of its pure inner forms. Here, $\mathcal{V}_{\mathrm{W},E}$ consists of tuples whose first component is an irreducible representation of a pure inner form of $\operatorname{U}(\mathrm{V})$, and whose second component is a character of a torus (of the aforementioned pure inner form) isomorphic to $L_{E}^{1}$. We move to explain our global result. Let $F=\mathbf{F}$ be a number field and let $K=\mathbf{K}$ be a quadratic field extension of $\mathbf{F}$. By choosing automorphic characters $\left(\chi_{\mathrm{W}},\chi_{\mathrm{V}}\right)$ of $\mathbb{A}_{\mathbf{K}}^{\times}$ we obtain a splitting $\operatorname{U}(\mathrm{V})\left(\mathbb{A}_{\mathbf{F}}\right)\times\operatorname{U}(\mathrm{W})\left(\mathbb{A}_{\mathbf{F}}\right)\to\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)\left(\mathbb{A}_{\mathbf{F}}\right)$. Let $i^{\prime}_{\mathrm{W}}\colon\operatorname{U}(\mathrm{W})\to L_{\mathbf{F}}^{1}$ be an admissible embedding, which in this case is an isomorphism. Suppose that $\mathbf{E}$ is an $n$-dimensional étale algebra over $\mathbf{F}$, such that there exists $\lambda\in\mathbf{E}^{\times}$ satisfying that $\mathrm{V}_{E,\lambda}$ is isomorphic to $\mathrm{V}$ as hermitian spaces, and such that $L_{\mathbf{E}}^{1}$ is anisotropic (equivalently, there is no embedding of $\mathbf{F}$-algebras $\mathbf{K}\hookrightarrow\mathbf{E}$). Let $\beta\colon\left[L_{\mathbf{F}}^{1}\right]\to\mathbb{C}^{\times}$ and $\alpha\colon\left[L_{\mathbf{E}}^{1}\right]\to\mathbb{C}^{\times}$ be automorphic characters. Our global result (Theorem 5.4) classifies when the $\alpha$-period is identically zero on the space of the global theta lift $\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$ of $\beta\circ i^{\prime}_{\mathrm{W}}$ from $\operatorname{U}(\mathrm{W})\left(\mathbb{A}_{\mathbf{F}}\right)$ to $\operatorname{U}(\mathrm{V})\left(\mathbb{A}_{\mathbf{F}}\right)$. As before, and as in the global Gan–Gross–Prasad conjectures, this classification is expressed in terms of root numbers and central values of $L$-functions.
@@
Let $i\colon L_{\mathbf{E}}^{1}\to\operatorname{U}(\mathrm{V})$ be an admissible embedding corresponding to $\lambda\in\mathbf{E}^{\times}$. The $\alpha\circ i^{-1}$-period $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ is non-zero on the global theta lift $\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$, that is,
$\int_{\left[L_{\mathbf{E}}^{1}\right]}f\left(i\left(t\right)\right)\overline{\alpha\left(t\right)}\mathrm{d}t\neq 0\text{ for some }f\in\Theta\left(\beta\circ i^{\prime}_{\mathrm{W}}\right),$
@@
    Character compatibility: $\beta=\left(\chi_{\mathrm{W}}^{-n}\cdot\chi_{\mathrm{V}}\right)\circ j_{\mathbb{A}_{\mathbf{F}}}^{-1}\cdot\alpha_{\restriction_{L_{\mathbf{F}}^{1}\left(\mathbb{A}_{\mathbf{F}}\right)}}.$
@@
    $\omega_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\left(\lambda\right)=\varepsilon_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\left(\alpha_{v,L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\cdot\chi_{\mathrm{W},v}^{-1}\circ N_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{K}\otimes_{\mathbf{F}}\mathbf{F}_{v}},\psi_{v},\delta\right).$
@@
    $\mathcal{L}\left(\alpha_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\cdot\chi_{\mathrm{W}}^{-1}\circ N_{L_{\mathbf{E}}/\mathbf{K}}\right)\neq 0.$
Here, $0\neq\delta\in\mathbf{K}^{\times}$ is a trace zero element, such that $\operatorname{disc}\mathrm{W}=\delta\left(\bmod N_{\mathbf{K}/\mathbf{F}}\left(\mathbf{K}^{\times}\right)\right)$. We refer the reader to Section 4.3 for the definition of the central $L$-function value $\mathcal{L}\left(\alpha_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\cdot\chi_{\mathrm{W}}^{-1}\circ N_{L_{\mathbf{E}}/\mathbf{K}}\right)$ and other notation appearing in the theorem. Notice that the first two conditions in Theorem 1.4 are equivalent to requiring that for every place $v$, the conditions in Theorem 1.2 hold. It is clear that the condition of the $\operatorname{Hom}$-space not vanishing for every $v$ is a necessary condition for the global period to not vanish. Moreover, the root number condition implies that the global root number equals one, i.e.,
$\prod_{v}\varepsilon_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\left(\alpha_{v,L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\cdot\chi_{\mathrm{W},v}^{-1}\circ N_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{K}\otimes_{\mathbf{F}}\mathbf{F}_{v}},\psi_{v},\delta\right)=1,$
and therefore the third condition regarding the central $L$-value $\mathcal{L}\left(\alpha_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\cdot\chi_{\mathrm{W}}^{-1}\circ N_{L_{\mathbf{E}}/\mathbf{K}}\right)$ can be satisfied. Similarly to the local case, we show in Corollary 5.6 that for characters $\alpha$ and $\beta$ satisfying the compatibility condition and such that the central $L$-function value in question does not vanish, there exists a unique non-degenerate hermitian space $\mathrm{H}$ of dimension $n$ and a unique class $i\in\Sigma_{\mathbf{E},\mathrm{H}}$, such that the period $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ does not identically vanish on $\Theta_{\mathrm{W},\mathrm{H}}\left(\beta\circ i^{\prime}_{\mathrm{W}}\right)$.

### v2

@@ Abstract.
We determine the restriction of Weil representations of unitary groups to maximal tori. In the local case, we show that the Weil representation contains a pair of compatible characters if and only if a root number condition holds. In the global case, we show that a torus period corresponding to a maximal anisotropic torus of the global theta lift of a character does not vanish if and only if the local condition is satisfied everywhere and a central value of an $L$-function does not vanish. Our proof makes use of the seesaw argument and of the well-known theta lifting results from $\operatorname{U}(1)$ to $\operatorname{U}(1)$. Our results are used in [1, 2] to construct Arthur packets for $G_{2}$.
Key words and phrases:
Theta correspondence, Periods, Special values of $L$-functions, Gan–Gross–Prasad conjectures
2020 Mathematics Subject Classification:
22E50, 11F27, 11F67
@@ 1. Introduction
Branching problems are a fascinating topic in representation theory and in the theory of automorphic representations. The most famous examples are the Gan–Gross–Prasad conjectures [9, 10, 11] and their refinements [18, 16, 33], extending the original conjectures of Gross–Prasad [13, 14].
In this paper, we study the restriction of Weil representations of unitary groups to maximal tori. Our results are both local and global, and they are similar in nature to the Gan–Gross–Prasad conjectures. Let us describe the problems we concern.
Let $F$ be a field with characteristic different than $2$ and let $K/F$ be a quadratic étale algebra with involution $x\mapsto x^{c}$, whose set of fixed points is $F$. Let $\mathrm{V}$ be a non-degenerate $n$-dimensional hermitian space over $K$, and let $\mathrm{W}$ be a non-degenerate one-dimensional skew-hermitian space over $K$.
When $F$ is a local field, we consider the following branching problem: given a maximal torus $\mathrm{T}$ of $\operatorname{U}(\mathrm{V})$ and characters $\alpha\colon\mathrm{T}\to\mathbb{C}^{\times}$ and $\beta\colon\operatorname{U}(\mathrm{W})\to\mathbb{C}^{\times}$, we would like to investigate whether the restriction of the Weil representation of the metaplectic group $\operatorname{Mp}_{\psi}\left(\mathrm{V},\mathrm{W}\right)$ to $\mathrm{T}\times\operatorname{U}(\mathrm{W})$ contains the representation $\alpha\boxtimes\beta$ as a sub-quotient. Reformulating this using the theta correspondence, this is equivalent to asking whether the space $\operatorname{Hom}_{\mathrm{T}}\left(\Theta\left(\beta\right),\alpha\right)$ is non-zero, where $\Theta\left(\beta\right)$ is the big theta lift of $\beta$ from $\operatorname{U}(\mathrm{W})$ to $\operatorname{U}(\mathrm{V})$.
Suppose that $F=\mathbf{F}$ is a number field and that $\mathbf{K}/\mathbf{F}$ is a quadratic field extension. For an algebraic group $G$, we write $\left[G\right]=G\left(\mathbf{F}\right)\backslash G\left(\mathbb{A}_{\mathbf{F}}\right)$. We consider the following branching problem: given a maximal torus $\mathrm{T}$ of $\operatorname{U}(\mathrm{V})$ and automorphic characters $\alpha\colon\left[\mathrm{T}\right]\to\mathbb{C}^{\times}$ and $\beta\colon\left[\operatorname{U}(\mathrm{W})\right]\to\mathbb{C}^{\times}$, we would like to investigate whether the $\alpha$-period of the global theta lift $\Theta\left(\beta\right)$ of $\beta$ from $\operatorname{U}(\mathrm{W})\left(\mathbb{A}_{\mathbf{F}}\right)$ to $\operatorname{U}(\mathrm{V})\left(\mathbb{A}_{\mathbf{F}}\right)$ is non-zero. That is, we are asking whether the assignment
@@
is identically zero on the space $\Theta\left(\beta\right)$. In order to avoid convergence issues, we will assume that the torus $\mathrm{T}$ is anisotropic in the global setting, so that the integrals in question converge absolutely.
Notice that when $\mathrm{V}$ is one-dimensional, the theta lift of $\beta$ is either zero or a character, and our problems reduce to determining whether $\Theta\left(\beta\right)$ equals $\alpha$ or not. This problem, of determining the theta lift from $\operatorname{U}(1)$ to $\operatorname{U}(1)$, is well understood. It dates back to Moen [24], Rogawski [28], and Harris–Kudla–Sweet [15] in the non-archimedean local field case, to Paul [25] in the archimedean local field case, to Minguez [22], Fang–Sun–Xue [5] and Gan [7] in the split local case, and to Rogawski [28], Yang [36] and Yamana [35] in the global case. See also Section 9 of [8] and the last paragraph of Section 7 of [8].
Our technique for solving these problems in the general case, where $\dim\mathrm{V}$ is arbitrary, involves a seesaw identity that reduces the problems to the well-known case discussed above. This idea has been used before by Gan and his collaborators, see for example [6, Sections 9 and 10], [7, Section 5] and [8, Section 10].
In order to state our results, we need a classification of maximal tori in $\operatorname{U}(\mathrm{V})$. In Section 2, we recall the classification given in [27]. Each maximal torus $\mathrm{T}\subset\operatorname{U}(\mathrm{V})$ corresponds to an étale algebra $E$ of degree $n$ over $F$ and an element $\lambda\in E^{\times}$, such that the space $\left(\mathrm{V}_{E,\lambda},\left\langle\cdot,\cdot\right\rangle_{\lambda}\right)$ is isomorphic to $\mathrm{V}$ as hermitian spaces, where $\mathrm{V}_{E,\lambda}=\operatorname{Res}_{L_{E}/K}L_{E}$, equipped with the hermitian product $\left\langle x,y\right\rangle_{\lambda}=\operatorname{tr}_{L_{E}/K}{\left(\lambda xy^{c}\right)}$, where $L_{E}=K\otimes_{F}E$. In this case, the maximal torus $\mathrm{T}_{E,\lambda}$ is isomorphic to the norm one torus of $L_{E}$, that is,
@@
We now move to describe our main results. In order to make the results look cleaner, we will no longer mention the one-dimensional skew-hermitian space $\mathrm{W}$ in the introduction, but instead use a trace zero element $\delta$ that encodes the discriminant of such space.
Suppose that $F$ is a local field. By choosing a trace zero element $\delta\in K^{\times}$, a character $\mu\colon K^{\times}\to\mathbb{C}^{\times}$ such that $\mu_{\restriction_{F^{\times}}}=\omega_{K/F}$ is the quadratic character given by local class field theory, and a non-trivial character $\psi\colon F\to\mathbb{C}^{\times}$, we can lift characters of $L_{F}^{1}$ to representations of $\operatorname{U}(\mathrm{V})$. Suppose that $E$ is an étale algebra of degree $n$ over $F$. Let $\beta\colon L_{F}^{1}\to\mathbb{C}^{\times}$, and $\alpha\colon L_{E}^{1}\to\mathbb{C}^{\times}$ be characters. We give the following answer (Theorem 5.1) to the local problem discussed above.
@@
$\operatorname{Hom}_{i\left(L_{E}^{1}\right)}\left(\Theta_{\delta,\mathrm{V},\mu,\psi}\left(\beta\right),\alpha\circ i^{-1}\right)$ (1)
@@
    Character compatibility: $\beta=\alpha_{\restriction_{L_{F}^{1}}}.$
@@
    Root number condition: $\omega_{L_{E}/E}\left(\lambda\right)=\varepsilon_{L_{E}/E}\left(\alpha_{L_{E}}\cdot\mu^{-1}\circ N_{L_{E}/K},\psi,\delta\right).$
@@
Here, $\Theta_{\delta,\mathrm{V},\mu,\psi}\left(\beta\right)$ is the big theta lift of $\beta$ to $\operatorname{U}(\mathrm{V})$ with respect to the data $\left(\delta,\mu,\psi\right)$, see Section 3.1.2. It is either zero or irreducible in our case (since $\beta$ is supercuspidal), and therefore equals the small theta lift $\theta_{\delta,\mathrm{V},\mu,\psi}\left(\beta\right)$. We refer the reader to Sections 2.2.2 and 3.3 for the definition of the vector of quadratic characters $\omega_{L_{E}/E}\left(\lambda\right)$, the vector of root numbers $\varepsilon_{L_{E}/E}\left(\alpha_{L_{E}}\cdot\mu^{-1}\circ N_{L_{E}/K},\psi,\delta\right)$, and other notation appearing in the theorem.
We also show that for characters $\alpha$ and $\beta$ satisfying the compatibility condition, there exists a unique non-degenerate hermitian space $\mathrm{H}$ of dimension $n$, up to isomorphism, and a unique admissible embedding $i\colon L_{E}^{1}\to\operatorname{U}(\mathrm{H})$, up to conjugation, such that the space (1) attached to $i$ is non-zero. More precisely, we show the following theorem (Corollary 5.4).
@@
$\sum_{\mathrm{H}\in\mathrm{Her}_{n}}\sum_{i\in\Sigma_{E,\mathrm{H}}}\dim_{\mathbb{C}}\operatorname{Hom}_{i\left(L_{E}^{1}\right)}\left(\Theta_{\delta,\mathrm{H},\mu,\psi}\left(\beta\right),\alpha\circ i^{-1}\right)=\begin{dcases}1&\beta=\alpha_{\restriction_{L_{F}^{1}}},\\ 0&\text{otherwise.}\end{dcases}$
@@
$\mathrm{Her}_{n}=\left\{\mathrm{H}\text{ is a non-degenerate hermitian space }\mid\dim\mathrm{H}=n\right\}/\text{isomorphism},$
and $i$ runs over representatives of classes of $\Sigma_{E,\mathrm{H}}$.
@@
$\mathcal{V}_{\delta,E}\left(\beta\right)=\bigcup_{\mathrm{H}\in\mathrm{Her}_{n}}\left\{\left(\Theta_{\delta,\mathrm{H},\mu,\psi}\left(\beta\right),\alpha\circ i^{-1}\right)\mid i\in\Sigma_{E,\mathrm{H}}\right\},$
is analogous to the local Vogan $L$-packet appearing in the Gan–Gross–Prasad conjectures, consisting of irreducible representations of the group and of its pure inner forms. Here, $\mathcal{V}_{\delta,E}$ consists of tuples whose first component is an irreducible representation of a pure inner form of $\operatorname{U}(\mathrm{V})$, and whose second component is a character of a torus (of the aforementioned pure inner form) isomorphic to $L_{E}^{1}$.
We move to explain our global result. Let $F=\mathbf{F}$ be a number field and let $K=\mathbf{K}$ be a quadratic field extension of $\mathbf{F}$.
By choosing a trace zero element $\delta\in\mathbf{K}^{\times}$, an automorphic character $\mu$ of $\mathbb{A}_{\mathbf{K}}^{\times}$, such that $\mu\restriction_{\mathbb{A}_{\mathbf{F}}^{\times}}=\omega_{{\mathbf{K}}/{\mathbf{F}}}$ is the quadratic character given by global class field theory, and a non-trivial character $\psi\colon\mathbf{F}\backslash\mathbb{A}\to\mathbb{C}^{\times}$, we can lift automorphic characters of $L_{\mathbf{F}}^{1}\left(\mathbb{A}_{\mathbf{F}}\right)$ to automorphic representations of $\operatorname{U}(\mathrm{V})\left(\mathbb{A}_{\mathbf{F}}\right)$. Suppose that $\mathbf{E}$ is an $n$-dimensional étale algebra over $\mathbf{F}$, such that there exists $\lambda\in\mathbf{E}^{\times}$ satisfying that $\mathrm{V}_{E,\lambda}$ is isomorphic to $\mathrm{V}$ as hermitian spaces, and such that $L_{\mathbf{E}}^{1}$ is anisotropic (equivalently, there is no embedding of $\mathbf{F}$-algebras $\mathbf{K}\hookrightarrow\mathbf{E}$). Let $\beta\colon\left[L_{\mathbf{F}}^{1}\right]\to\mathbb{C}^{\times}$ and $\alpha\colon\left[L_{\mathbf{E}}^{1}\right]\to\mathbb{C}^{\times}$ be automorphic characters. Our global result (Theorem 5.6) classifies when the $\alpha$-period is identically zero on the space of the global theta lift $\Theta_{\delta,\mathrm{V},\mu,\psi}\left(\beta\right)$ of $\beta$ to $\operatorname{U}(\mathrm{V})\left(\mathbb{A}_{\mathbf{F}}\right)$. As before, and as in the global Gan–Gross–Prasad conjectures, this classification is expressed in terms of root numbers and central values of $L$-functions.
@@
Let $i\colon L_{\mathbf{E}}^{1}\to\operatorname{U}(\mathrm{V})$ be an admissible embedding corresponding to $\lambda\in\mathbf{E}^{\times}$. The $\alpha\circ i^{-1}$-period $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ is non-zero on the global theta lift $\Theta_{\delta,\mathrm{V},\mu,\psi}\left(\beta\right)$, that is,
$\int_{\left[L_{\mathbf{E}}^{1}\right]}f\left(i\left(t\right)\right)\overline{\alpha\left(t\right)}\mathrm{d}t\neq 0\text{ for some }f\in\Theta_{\delta,\mathrm{V},\mu,\psi}\left(\beta\right),$
@@
    Character compatibility: $\beta=\alpha_{\restriction_{L_{\mathbf{F}}^{1}\left(\mathbb{A}_{\mathbf{F}}\right)}}.$
@@
    $\omega_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\left(\lambda\right)=\varepsilon_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\left(\alpha_{v,L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\cdot\mu_{v}^{-1}\circ N_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{K}\otimes_{\mathbf{F}}\mathbf{F}_{v}},\psi_{v},\delta\right).$
@@
    $\mathcal{L}\left(\alpha_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\cdot\mu^{-1}\circ N_{L_{\mathbf{E}}/\mathbf{K}}\right)\neq 0.$
We refer the reader to Section 4.3 for the definition of the central $L$-function value $\mathcal{L}\left(\alpha_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\cdot\mu^{-1}\circ N_{L_{\mathbf{E}}/\mathbf{K}}\right)$ and other notation appearing in the theorem.
Notice that the first two conditions in Theorem 1.4 are equivalent to requiring that for every place $v$, the conditions in Theorem 1.2 hold. It is clear that the condition of the $\operatorname{Hom}$-space not vanishing for every $v$ is a necessary condition for the global period to not vanish. Moreover, the root number condition implies that the global root number equals one, i.e.,
$\prod_{v}\varepsilon_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{E}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\left(\alpha_{v,L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}}\cdot\mu_{v}^{-1}\circ N_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbf{F}_{v}/\mathbf{K}\otimes_{\mathbf{F}}\mathbf{F}_{v}},\psi_{v},\delta\right)=1,$
and therefore the third condition regarding the central $L$-value $\mathcal{L}\left(\alpha_{L_{\mathbf{E}}\otimes_{\mathbf{F}}\mathbb{A}_{\mathbf{F}}}\cdot\mu^{-1}\circ N_{L_{\mathbf{E}}/\mathbf{K}}\right)$ can be satisfied.
Similarly to the local case, we show in Corollary 5.8 that for characters $\alpha$ and $\beta$ satisfying the compatibility condition and such that the central $L$-function value in question does not vanish, there exists a unique non-degenerate hermitian space $\mathrm{H}$ of dimension $n$ and a unique class $i\in\Sigma_{\mathbf{E},\mathrm{H}}$, such that the period $\mathcal{P}_{i\left(L_{\mathbf{E}}^{1}\right),\alpha\circ i^{-1}}$ does not identically vanish on $\Theta_{\delta,\mathrm{H},\mu,\psi}\left(\beta\right)$.